# Quantum-limited optical delay sensing across an enhanced dynamic range by frequency-resolving two-photon interference

### v1

@@ Acknowledgements
$\begin{split}P_{\eta}(W,\Delta\omega,X)&=\frac{\gamma^{2}}{2\pi\sigma^{2}}\exp\left[-\frac{(W-\Omega_{0})^{2}}{\sigma^{2}}\right]\exp\left[-\frac{\Delta\omega^{2}}{4\sigma^{2}}\right]\frac{1+\alpha(X)\eta^{2}\cos\left(\Delta\omega\Delta t\right)}{2}\\ &=G_{\Omega_{0},\sigma^{2}}(W)P_{\eta}(\Delta\omega,X),\end{split}$ (SM7)
@@
$P_{\eta}(\Delta\omega,X)=\int dWP_{\eta}(W,\Delta\omega,X)=\gamma^{2}C(\Delta\omega)\frac{1+\alpha(X)\eta^{2}\cos\left(\Delta\omega\Delta t\right)}{2},$ (SM8)
@@
$C(\Delta\omega)=\frac{1}{\sqrt{4\pi\sigma^{2}}}\exp\left[-\frac{\Delta\omega^{2}}{4\sigma^{2}}\right].$ (SM9)
@@
$\begin{split}\sum_{X=B,C}\int dWd\omega\frac{(\partial_{\Delta t}P_{\eta}(W,\omega,X))^{2}}{P_{\eta}(W,\omega,X)}&=\sum_{X=B,C}\int dWd\omega\frac{(G_{\Omega_{0},\sigma^{2}}(W)\partial_{\Delta t}P_{\eta}(\omega,X))^{2}}{G_{\Omega_{0},\sigma^{2}}(W)P_{\eta}(\omega,X)}\\ &=\sum_{X=B,C}\int dWG_{\Omega_{0},\sigma^{2}}(W)\int d\omega\frac{(\partial_{\Delta t}P_{\eta}(\omega,X))^{2}}{P_{\eta}(\omega,X)}\\ &=\sum_{X=B,C}\int d\omega\frac{(\partial_{\Delta t}P_{\eta}(\omega,X))^{2}}{P_{\eta}(\omega,X)}.\end{split}$ (SM10)
SM0.10 Finite resolution probabilities
@@
$P\bigl(X,\tau,\omega,\sigma,\eta\bigr)=\frac{1}{\sqrt{4\pi\sigma^{2}}}\exp\!\biggl(-\frac{\omega^{2}}{4\sigma^{2}}\biggr)\,\frac{1+\alpha(X)\,\eta\,\cos(\tau\,\omega)}{2}\,.$ (SM11)
@@
$\begin{split}P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)&=\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega\\ &=\frac{1}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\mathrm{e}^{-\varpi^{2}}\frac{1+\alpha(X)\eta\cos(2\sigma\tau\varpi)}{2}d\varpi,\end{split}$ (SM12)
@@
$\begin{split}P_{n>0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)&=\int_{(n-1)\epsilon}^{(n+1)\epsilon}f_{n}\left(\frac{\omega}{\epsilon}\right)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega\\ &=\frac{1}{\sqrt{\pi}}\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}\frac{1+\alpha(X)\eta\cos(2\sigma\tau\varpi)}{2}d\varpi\end{split}$ (SM13)
@@
$f_{n}\left(\frac{\omega}{\epsilon}\right)=\begin{cases}\frac{\omega}{\epsilon}-n+1\qquad\mathrm{for}\qquad(n-1)\epsilon\leq\omega<n\epsilon\\ n+1-\frac{\omega}{\epsilon}\qquad\mathrm{for}\qquad n\epsilon\leq\omega\leq(n+1)\epsilon\end{cases}\qquad.$ (SM14)
@@
$F_{n}(\tau)=\sum_{X=A,B}\frac{1}{P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)}\left(\frac{\partial P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)}{\partial\tau}\right)^{2},$ (SM16)
@@
$\begin{split}F_{0}(\tau)&=\sum_{X=A,B}\frac{1}{P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)}\left(\frac{\partial P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)}{\partial\tau}\right)^{2}\\ &=\sum_{X=A,B}\frac{1}{\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega}\left(\frac{\partial}{\partial\tau}\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega\right)^{2}\end{split}$ (SM17)
@@
$\begin{split}\left(\frac{\partial}{\partial\tau}\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega\right)^{2}&=\left(\frac{\sigma\eta}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}.\end{split}$ (SM18)
@@
$\begin{split}F_{0}(\tau)&=\left(\frac{\sigma\eta}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\sum_{X=A,B}\frac{1}{\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega}\\ &=\frac{\left(\frac{\sigma\eta}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\left(\frac{1}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)}{\left(\frac{1}{2\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)^{2}-\left(\frac{\eta}{2\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\cos(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}}\\ &=\frac{4\sigma^{2}\eta^{2}}{\sqrt{\pi}}\frac{\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)}{\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)^{2}-\left(\eta\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\cos(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}}.\end{split}$ (SM19)
@@
$F_{0}(\tau)=H\frac{2\eta^{2}I_{0,\mathrm{num}}}{\sqrt{\pi}I_{0,\mathrm{den}}},$ (SM20)
@@
$I_{0,\mathrm{num}}=\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)$ (SM21)
@@
$I_{0,\mathrm{den}}=\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)^{2}-\left(\eta\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\cos(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}.$ (SM22)
@@
$F_{n>0}(\tau)=H\frac{2\eta^{2}I_{n>0,\mathrm{num}}}{\sqrt{\pi}I_{n>0,\mathrm{den}}},$ (SM23)
@@
$I_{n>0,\mathrm{num}}=\left(\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\left(\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)$ (SM24)
@@
$I_{n>0,\mathrm{den}}=\left(\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)^{2}-\left(\eta\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}\cos(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}.$ (SM25)
Figure SM1: Experimental layout with two sources. A 775 nm pulsed Ti:Sapphire laser with 1.3 ps duration and 80 MHz repetition rate pumps two photon pair sources. The sources comprise of a aperiodicity-poled KTP crystal housed in a Sagnac configuration, generating spectrally degenerate Gaussian profile photon pairs at 1550 nm. One photon from each pair is separated using dichroic polarising beam splitters (dPBS) and coupled to a single-mode fiber (SMF) to be detected imediately on superconducting nanowire detectors (SNSPD) and time-stamped on Hydra harp time tagger. The other pair from each source is then coupled to SMF and launched into free space onto a 50:50 beam splitter. The polarisation of each independent photon is prepared using a linear plate polariser set to Horizontal to match the polarisation degree of freedom. The Spectral component is also matched by calibrating the temperature of the crystal ovens to yield the highest independent-HOM visibility, found to be 91%. Each path exiting the beam splitter is coupled to 50km SMF to induce chromatic dispersion and delay the photon packets in time. A time-of-flight spectrometer converts the time difference between the arrival of a photon and the arrival of the Ti:Sapphire sync pulse to estimate the frequency component after experience dispersion. Finally each path is multiplexed into two spatial modes to act as a pseudo number-resolving detectors.
Table 1: Experimental parameters
Metric Gaussian $\sigma$ FWHM
Coherence length $\tau_{c}$ 1.24 2.92
Temporal bandwidth $\tau$ [ps] 0.98 2.31
Spectral bandwidth $\sigma/2\pi$ [GHz] 81 190
Figure SM3: The Allan variance of the 2-fold samples is plotted to assess the stability of the estimates over a long period of time.
[…]

### v2

@@ Acknowledgements
$\begin{split}P_{\eta}(W,\Delta\omega,X)&=\frac{\gamma^{2}}{2\pi\sigma^{2}}\exp\left[-\frac{(W-\Omega_{0})^{2}}{\sigma^{2}}\right]\exp\left[-\frac{\Delta\omega^{2}}{4\sigma^{2}}\right]\frac{1+\alpha(X)\eta^{2}\cos\left(\Delta\omega\Delta t\right)}{2}\\ &=G_{\Omega_{0},\sigma^{2}}(W)P_{\eta}(\Delta\omega,X),\end{split}$ (SM8)
@@
$P_{\eta}(\Delta\omega,X)=\int dWP_{\eta}(W,\Delta\omega,X)=\gamma^{2}C(\Delta\omega)\frac{1+\alpha(X)\eta^{2}\cos\left(\Delta\omega\Delta t\right)}{2},$ (SM9)
@@
$C(\Delta\omega)=\frac{1}{\sqrt{4\pi\sigma^{2}}}\exp\left[-\frac{\Delta\omega^{2}}{4\sigma^{2}}\right].$ (SM10)
@@
$\begin{split}\sum_{X=B,C}\int dWd\omega\frac{(\partial_{\Delta t}P_{\eta}(W,\omega,X))^{2}}{P_{\eta}(W,\omega,X)}&=\sum_{X=B,C}\int dWd\omega\frac{(G_{\Omega_{0},\sigma^{2}}(W)\partial_{\Delta t}P_{\eta}(\omega,X))^{2}}{G_{\Omega_{0},\sigma^{2}}(W)P_{\eta}(\omega,X)}\\ &=\sum_{X=B,C}\int dWG_{\Omega_{0},\sigma^{2}}(W)\int d\omega\frac{(\partial_{\Delta t}P_{\eta}(\omega,X))^{2}}{P_{\eta}(\omega,X)}\\ &=\sum_{X=B,C}\int d\omega\frac{(\partial_{\Delta t}P_{\eta}(\omega,X))^{2}}{P_{\eta}(\omega,X)}.\end{split}$ (SM11)
SM10 Finite resolution probabilities
@@
$P\bigl(X,\tau,\omega,\sigma,\eta\bigr)=\frac{1}{\sqrt{4\pi\sigma^{2}}}\exp\!\biggl(-\frac{\omega^{2}}{4\sigma^{2}}\biggr)\,\frac{1+\alpha(X)\,\eta\,\cos(\tau\,\omega)}{2}\,.$ (SM12)
@@
$\begin{split}P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)&=\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega\\ &=\frac{1}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\mathrm{e}^{-\varpi^{2}}\frac{1+\alpha(X)\eta\cos(2\sigma\tau\varpi)}{2}d\varpi,\end{split}$ (SM13)
@@
$\begin{split}P_{n>0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)&=\int_{(n-1)\epsilon}^{(n+1)\epsilon}f_{n}\left(\frac{\omega}{\epsilon}\right)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega\\ &=\frac{1}{\sqrt{\pi}}\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}\frac{1+\alpha(X)\eta\cos(2\sigma\tau\varpi)}{2}d\varpi\end{split}$ (SM14)
@@
$f_{n}\left(\frac{\omega}{\epsilon}\right)=\begin{cases}\frac{\omega}{\epsilon}-n+1\qquad\mathrm{for}\qquad(n-1)\epsilon\leq\omega<n\epsilon\\ n+1-\frac{\omega}{\epsilon}\qquad\mathrm{for}\qquad n\epsilon\leq\omega\leq(n+1)\epsilon\end{cases}\qquad.$ (SM15)
@@
$F_{n}(\tau)=\sum_{X=A,B}\frac{1}{P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)}\left(\frac{\partial P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)}{\partial\tau}\right)^{2},$ (SM17)
@@
$\begin{split}F_{0}(\tau)&=\sum_{X=A,B}\frac{1}{P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)}\left(\frac{\partial P_{0}\bigl(X,\tau,\epsilon,\sigma,\eta\bigr)}{\partial\tau}\right)^{2}\\ &=\sum_{X=A,B}\frac{1}{\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega}\left(\frac{\partial}{\partial\tau}\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega\right)^{2}\end{split}$ (SM18)
@@
$\begin{split}\left(\frac{\partial}{\partial\tau}\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega\right)^{2}&=\left(\frac{\sigma\eta}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}.\end{split}$ (SM19)
@@
$\begin{split}F_{0}(\tau)&=\left(\frac{\sigma\eta}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\sum_{X=A,B}\frac{1}{\int_{0}^{\epsilon}\Bigl(1-\tfrac{\omega}{\epsilon}\Bigr)\,P\bigl(X,\tau,\omega,\sigma,\eta\bigr)\,\mathrm{d}\omega}\\ &=\frac{\left(\frac{\sigma\eta}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\left(\frac{1}{\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)}{\left(\frac{1}{2\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)^{2}-\left(\frac{\eta}{2\sqrt{\pi}}\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\cos(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}}\\ &=\frac{4\sigma^{2}\eta^{2}}{\sqrt{\pi}}\frac{\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)}{\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)^{2}-\left(\eta\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\cos(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}}.\end{split}$ (SM20)
@@
$F_{0}(\tau)=H\frac{2\eta^{2}I_{0,\mathrm{num}}}{\sqrt{\pi}I_{0,\mathrm{den}}},$ (SM21)
@@
$I_{0,\mathrm{num}}=\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)$ (SM22)
@@
$I_{0,\mathrm{den}}=\left(\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)^{2}-\left(\eta\int_{0}^{\epsilon^{\prime}}\Bigl(1-\tfrac{\varpi}{\epsilon^{\prime}}\Bigr)\,\mathrm{e}^{-\varpi^{2}}\cos(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}.$ (SM23)
@@
$F_{n>0}(\tau)=H\frac{2\eta^{2}I_{n>0,\mathrm{num}}}{\sqrt{\pi}I_{n>0,\mathrm{den}}},$ (SM24)
@@
$I_{n>0,\mathrm{num}}=\left(\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}\varpi\sin(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}\left(\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)$ (SM25)
@@
$I_{n>0,\mathrm{den}}=\left(\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}d\varpi\right)^{2}-\left(\eta\int_{(n-1)\epsilon^{\prime}}^{(n+1)\epsilon^{\prime}}f_{n}\left(\frac{\varpi}{\epsilon^{\prime}}\right)\,\mathrm{e}^{-\varpi^{2}}\cos(2\sigma\tau\varpi)\,\mathrm{d}\varpi\right)^{2}.$ (SM26)
Figure SM4: The Allan variance of the 2-fold samples is plotted to assess the stability of the estimates over a long period of time.